If the error function for 4 experimental points equals to 50.4, then the root mean square (rms) for these points equals to


1
Expert's answer
2021-06-23T06:57:55-0400

Solution:

Given, Error function, "E=50.4"

And number of observations, "N=4"

Then, "RMS=\sqrt{\dfrac{E^2}{N}}=\sqrt{\dfrac{50.4^2}{4}}=25.2"